# Thunderscapes: Simulating the Dynamics of Mesoscale Convective System

### v3

@@ 7. VALIDATION
We adopt the evaluation methods outlined in (Shen et al., 2020) and (Cesana et al., 2019), utilizing altitude-based quantitative data to analyze cloud behavior across different atmospheric layers. The results are presented in Figure 8, which illustrates the relationship between cloud fraction and storm height. Notably, the maximum cloud fraction corresponds to the storm’s peak altitude, consistent with the classical vertical distribution of cumulonimbus clouds in meteorology.
Figure 8. Comparative visualization of cloud fractions under different storm systems. From left to right: single-cell, multi-cell, squall line, and supercell. Each panel highlights the cloud fraction structure typical of these systems.
For further validation, we specifically selected four regions frequently studied in thunderstorm meteorology (Uman, 2001; Harris and Carvalho, 2018). Following the methodology from (Herrera et al., 2021), we examine the temporal evolution of the thunderstorm described in section 6.2. The evolution of cloud coverage is compared with real-world weather data from public sources, showing dynamic changes in the distribution of cloud cover. This comparison helps assess the accuracy of the model in capturing cloud coverage patterns observed in nature.we present the maximum storm development height along with the corresponding cloud fraction, providing a detailed overview of the storm’s vertical structure. This information is used to validate the model’s ability to simulate the typical altitude and distribution of clouds during storm development.
In line with the evaluation method in (Formenton et al., 2013), we analyze the evolution of lightning flash rates to capture the electrical activity of thunderstorms. The flash rate increases as the storm develops and gradually declines to zero as the storm dissipates. This trend reflects the typical lifecycle of a thunderstorm and its charge dynamics. Additionally, the frequency of lightning occurrences is examined in relation to altitude.
Figure 9 provides a comparison across Florida, New Mexico, Japan, and California, illustrating the model’s ability to capture a wide range of climatic behaviors and atmospheric dynamics, and confirming its consistency with observed storm structures and lightning patterns.
@@ 8. CONCLUSION AND FUTURE WORK
We have developed a comprehensive framework for simulating atmospheric phenomena during thunderstorms, focusing on thundercloud development, dissipation, and lightning generation through cloud electrification processes. This framework integrates a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processes to ensure the consistent coupling of cloud formation and lightning dynamics. The model simulates key atmospheric quantities, including microphysical properties such as vapor, cloud water, ice, and precipitated rain and snow, as well as electrodynamic properties like static charge distribution and lightning flashes. Our framework reproduces various thundercloud types, including single-cell, multicell, squall line, supercell, and atmospheric river formations, and simulates lightning dynamics using the Dielectric Breakdown Model (DBM), capturing both CG and IC lightning. Validation against real-world weather data demonstrates the model’s accuracy in simulating cloud structure, temperature evolution, cloud coverage, relative humidity, and lightning flash rates, with results benchmarked against national weather services and observed lightning activity.
Figure 9. Comparison of cloud coverage evolution (top row) ,cloud fraction (middle row) and lightning activity evolution(bottom row)across four distinct regions: Florida, New Mexico, Japan, and California. The figures showcase the distinct patterns of atmospheric dynamics influenced by regional environmental conditions.
Future work will focus on expanding the framework to incorporate advanced atmospheric concepts and support a broader range of thunderstorm and lightning phenomena. Enhancements to thunderstorm microphysics, such as more detailed electrification processes and turbulence modeling, will improve the realism of lightning generation. The framework will also be extended to simulate additional thunderstorm types, including Mesoscale Convective Complex (MCC), Mesoscale Convective Vortex (MCV), and derechos, broadening its coverage of mesoscale systems. Furthermore, new lightning types such as anvil crawlers, bolts from the blue, and sheet lightning will be introduced, requiring refined electric field and branching models. Finally, scalability and real-time performance will be improved by leveraging advanced numerical techniques and hardware optimizations, enabling more detailed and efficient simulations.

### v4

@@ 7. VALIDATION
The validation of our model involves presenting spatially simulated thunderstorm structures alongside meteorological characteristics and comparing the temporally simulated results with real-world weather data. This includes analyzing cloud fraction profiles to assess cloud formation and structure, tracking the evolution of cloud coverage against real-life data from national weather services [^3] https://www.visualcrossing.com/weather/weather-data-services/, and evaluating the temporal variation in lightning flash rates.
Figure 5. Visualization comparing cloud fraction models and simulation results for four storm types: single cell, multi-cell, squall line, and supercell. Top row: Cloud fraction models, where orange represents our method, and blue represents the method of Hädrich et al. (Hädrich et al., 2020). Middle row: Results generated using Hädrich et al. (Hädrich et al., 2020). Bottom row: Results generated by our method. The visualization highlights differences in cloud fraction structures and thunderstorm characteristics captured by the two approaches.
7.1. Thunderstorm structure
We adopt the evaluation methods outlined in (Shen et al., 2020) and (Cesana et al., 2019), utilizing altitude-based quantitative data to analyze cloud behavior across various atmospheric layers. To ensure a fair comparison, we evaluate our framework alongside the cumulonimbus simulation model Stormscapes (Hädrich et al., 2020) using identical input fields and consistent initial conditions.
The results, shown in Figure 5, illustrate the relationship between cloud fraction and the structure of thunderstorms at their maximum development height as modeled by the two approaches. While both models perform well for simpler thunderstorm types, such as single cell thunderstorms, notable differences arise for more complex systems, including multi-cell thunderstorms, squall lines, and supercells. At higher altitudes, the Stormscapes struggles to capture the expected extension of cloud coverage, particularly the anvil cloud, which is a critical feature of complex thunderstorms. The anvil, characterized by its solid appearance, sharp, well-defined edges, and downwind spread influenced by upper-level steering winds, is not accurately represented in Stormscapes.
This limitation arises because the Stormscapes relies on the classical Kessler-style precipitation parameterization, which only accounts for phase transitions among vapor, rain, and cloud water. In contrast, our framework employs a Grabowski-style microphysics scheme, which incorporates a finer classification of phase transitions, including vapor, rain, snow, cloud water, and ice. This enhanced granularity enables more accurate modeling of atmospheric buoyancy and better aligns with meteorological observations, making our model particularly well-suited for realistic MCS simulations.Moreover, the Stormscapes model does not account for hydrometeor electrification processes, limiting its ability to simulate lightning dynamics during thunderstorm development and to reproduce consistent atmospheric phenomena.
Figure 6. Comparison of MCS across four distinct regions: Florida, New Mexico, Japan, and California. Top row: Cloud coverage evolution over a 24-hour period during thunderstorm events. Middle row: Cloud fraction analysis, illustrating the structure and distribution of clouds at the maximum development height of thunderstorms. Bottom row: Lightning activity evolution, capturing the variations in lightning activity as thunderstorms develop. The figures highlight the unique patterns shaped by regional environmental conditions.
7.2. Consistent Thunderstorm System
We proceed to analyze and quantify the simulation data discussed in Section 6.2.Figure 6 provides a comparative overview of MCS characteristics across four regions: Florida, New Mexico, Japan, and California. Additionally, Figure 7 illustrates the development of MCS over specific evolution times, offering insights into temporal dynamics.
Our analysis begins by comparing the simulated cloud coverage attributes with 24-hour observational meteorological data. Next, we assess the cloud fraction at the maximum development height of thunderstorms in each region, providing critical insights into the vertical structure and dynamics of the thunderstorm systems.
Using the evaluation methodology described in (Formenton et al., 2013), we further analyze the evolution of lightning flash rates to capture the electrical activity within the thunderstorms. The flash rate follows a characteristic trend: increasing during the thunderstorm’s intensification phase, peaking at maturity, and gradually declining as the system dissipates. Additionally, we examine the altitude of lightning generation, revealing the vertical distribution of electrical activity. The alignment of lightning flash rates with the spatial and temporal evolution of the thunderstorm systems demonstrates the model’s consistency in realistically simulating the life cycle of MCS.
Figure 7. Different regions’ MCS evolution over time. First row: Florida at times $t_{1}=6.56\text{h},t_{2}=9.58\text{h},t_{3}=13.44\text{h},t_{4}=20.51\text{h}$. Second row: New Mexico at times $t_{1}=6.12\text{h},t_{2}=11.45\text{h},t_{3}=17.70\text{h},t_{4}=22.32\text{h}$. Third row: Japan at times $t_{1}=5.36\text{h},t_{2}=9.34\text{h},t_{3}=10.36\text{h},t_{4}=10.38\text{h}$. Fourth row: California at times $t_{1}=9.89\text{h},t_{2}=11.04\text{h},t_{3}=11.80\text{h},t_{4}=16.26\text{h}$.
@@ 8. CONCLUSION AND FUTURE WORK
We have developed a physically based simulation framework for visually realistic MCS. This framework integrates a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processes to ensure the consistent coupling of thunderstorms formation and lightning dynamics. Our framework reproduces various thunderstorms types within a MCS.Validation against real-world weather data demonstrates the model’s consistency in simulating cloud structure, cloud coverage, and lightning flash rates, with results benchmarked against national weather services.
Future work will focus on expanding the framework to incorporate advanced atmospheric concepts and support a broader range of thunderstorm and lightning phenomena. Enhancements to thunderstorm microphysics, such as more detailed electrification processes and turbulence modeling, will improve the realism of lightning generation. The framework will also be extended to simulate additional thunderstorm types, including Mesoscale Convective Complex (MCC), Mesoscale Convective Vortex (MCV), and derechos, broadening its coverage of mesoscale systems. Furthermore, new lightning types such as anvil crawlers, bolts from the blue, and sheet lightning will be introduced, requiring refined electric field and branching models. Finally, scalability and performance will be enhanced by leveraging advanced numerical techniques and improving hardware utilization. Currently, our numerical method employs a uniform grid.Transitioning to an adaptive grid approach (Raateland et al., 2022) appears to be a promising direction for achieving higher performance.